OpenStudy (anonymous):

Choose the equation of the line passing through the point (3, -2) and parallel to y = 3x - 6. y = 3x + 7 y = 3x - 11 y = -3x + 11 y = -3x - 12 My answer is y = 3x - 11 Am I right? :)

OpenStudy (anonymous):

Parallel lines have equal gradients. The gradient of the first line is 3 because it's 3x. Therefore the gradient of the other line must also be 3. Then use the formula for the equation of a line form there.\[y-y1 = m(x - x1)\]So you get y - (-2) = 3 (x - 3) which expands to be y + 2 = 3x - 9. And if you move the 2 across then you get y=3x-11 so yes you are correct sir!

OpenStudy (anonymous):

Thank you so much! :)